# ARise: Towards Knowledge-Augmented Reasoning via Risk-Adaptive Search

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities and are receiving increasing attention to enhance their reasoning through scaling test-time compute. However, their application in open-ended, knowledge-intensive, complex reasoning scenarios is still limited. Reasoning-oriented methods struggle to generalize to open-ended scenarios due to implicit assumptions of complete world knowledge. Meanwhile, knowledge-augmented reasoning (KAR) methods fails to address two core challenges: 1) error propagation, where errors in early steps cascade through the chain, and 2) verification bottleneck, where the explore–exploit trade-off arises in multi-branch decision processes. To overcome these limitations, we introduce ARise, a novel framework that integrates risk assessment of intermediate reasoning states with dynamic retrieval-augmented generation (RAG) within a Monte Carlo tree search paradigm. This approach enables effective construction and optimization of reasoning plans across multiple maintained hypothesis branches. Experimental results show that ARise significantly outperforms the state-of-the-art KAR methods by up to 23.10%, and the latest RAG-equipped large reasoning models by up to 25.37%. Our project page is at https://opencausalab.github.io/ARise.
@@
Yize Zhang1,2,3thanks: Equal contribution. Tianshu Wang4,5,711footnotemark: 1 Sirui Chen1,6 Kun Wang4 Xingyu Zeng4 Hongyu Lin5 Xianpei Han5 Le Sun5 Chaochao Lu1,2thanks: Corresponding author. 1Shanghai AI Laboratory  2Shanghai Innovation Institute  3Shanghai Jiao Tong University 4SenseTime  5Institute of Software, Chinese Academy of Sciences  6Tongji University 7Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences ez220523@sjtu.edu.cn, tianshu2020@iscas.ac.cn, luchaochao@pjlab.org.cn
@@ 1 Introduction
Figure 1: Error propagation and verification bottleneck. Prior works on knowledge-augmented reasoning fails to address two core challenges: 1) error propagation, where errors in early steps cascade through the chain, and 2) verification bottleneck, where the explore–exploit trade-off arises in multi-branch decision processes.
@@
Meanwhile, current knowledge-augmented reasoning (KAR) methods, as illustrated in Figure 1, are hindered by error propagation and verification bottleneck, which undermine reasoning reliability. To acquire knowledge for reasoning, retrieval-augmented generation (RAG) has shown as an effective way by dynamically retrieving documents as intermediate results Lewis et al. (2020); Liu et al. (2024). Prompt-based methods further extend KAR through chain-of-thought (CoT) prompting, which decomposes complex reasoning into sub-steps and iteratively retrieves relevant knowledge as reasoning proceeds Zhao et al. (2023a); Yu et al. (2024b); Li et al. (2024). However, this approach are plagued by error propagation, where errors in early steps can cascade through the chain. While search-based methods can mitigate error propagation through maintaining multiple hypothesis branches, verification bottleneck limits the effective explore–exploit tradeoff in KAR. Existing verification solutions remain unsatisfactory as they rely on error-prone self-verification Stechly et al. (2024); Wang et al. (2023b); Zhang et al. (2024b), or on specific verifier training Setlur et al. (2024); Zhang et al. (2024a).
To overcome these limitations, we present a novel framework, ARise, towards knowledge-Augmented Reasoning via risk-adaptive search. As shown in Figure 2, ARise consists of three components: reasoning state generation, Monte Carlo tree search (MCTS), and risk assessment. Specifically, ARise iterative refines reasoning steps through decomposition, retrieval-then-reasoning to provide fine-grained knowledge for LLMs (§ 2.1). MCTS treats each step as a node in the search tree, expanding linear reasoning to mitigate error propagation by enabling focused exploration of promising reasoning state and allowing backtracking when necessary (§ 2.2). Risk assessment leverages Bayesian risk minimization to evaluate the uncertainty of each state, dynamically balancing explore–exploit trade-off to guide the search towards both reliable and novel reasoning directions (§ 2.3). In this way, ARise enables robust and efficient complex reasoning by combining structured decomposition, knowledge retrieval, and risk-adaptive exploration in a unified framework.
We conducted comprehensive experiments with multiple LLMs on three challenging multi-hop question answering (QA) benchmarks that require complex reasoning and knowledge integration. Experimental results demonstrate that ARise significantly outperforms the state-of-the-art (SOTA) KAR methods, with an average of 23.10% and 15.52% improvement in accuracy and F1. In addition, when compared to latest LRMs DeepSeek-AI et al. (2025) equipped with RAG, ARise also improve the average accuracy and F1 of 4.04% and 25.37%. These results verify the effectiveness of ARise for open-ended, knowledge-intensive, complex reasoning tasks.
@@
  • We propose a knowledge-augmented framework for open-ended complex reasoning and design a risk-adaptive MCTS algorithm to balance explore–exploit tradeoff for reasoning.
  • We conduct comprehensive experiments to verify the effectiveness of ARise and to demonstrate it outperforms SOTA KAR methods and lastest LRMs equipped with RAG.
  • We provide empirical insights that 1) search-based wide reasoning can explore more solutions than learning-based deep reasoning, and 2) ARise progressively approach optimal performance through model size scaling.
@@ 2 The ARise Method
Figure 2: Pipeline of ARise. ARise iterative refines reasoning steps through decomposition, retrieval-then-reasoning, providing fine-grained knowledge for LLMs (§ 2.1). MCTS treats each step as a node in the search tree, expanding linear reasoning to mitigate error propagation by enabling exploration of reasoning paths and allowing backtracking when necessary (§ 2.2). Risk assessment leverages Bayesian risk minimization to evaluate the quality of each state, dynamically optimizing action strategies to guide the search towards promising directions (§ 2.3).
@@
  • Monte Carlo Tree Search: MCTS transforms a sequence of interconnected nodes into a tree structure. Each node can undergo a simulated rollout under guidance. The local value incorporating future reward is able to be updated without incurring the cost of taking an actual forward step.
  • Risk Assessment: We designed the Risk-Value function to risk-assess the intermediate reasoning states at each node. The policy model is capable of dynamically formulating and adjusting action strategies based on the actual risk associated with each branch.
@@ 2.1 Reasoning State Generation
To define steps more clearly and granularly, we prompt LLMs to perform problem decomposition and retrieval-then-reasoning in an interleaved manner. The two consecutive actions together constitute a single step. Intermediate results at each step are continuously appended to the entire reasoning state and serve as new inputs for subsequent steps, progressively approaching the final solution to complex tasks. This approach, where intermediate sub-tasks and their labels are concatenated to the original task’s input to form a new input sequence, is widely applied in compounded tasks (Wies et al., 2023; Rajani et al., 2019; Cobbe et al., 2021). Specifically, at the $i^{\text{th}}$ step [^2] We use the bold notation for vectors and non-bold notation for scalars., the input comprises the original problem $\mathbf{q}$ and the intermediate results ${\mathbf{r_{1}},\mathbf{r_{2}},\mathbf{\dots},\mathbf{r_{i-1}}}$ from previous steps, with the latter forming the reasoning state $\mathbf{s_{i-1}}=\mathbf{r_{1}}\oplus\mathbf{r_{2}}\oplus\mathbf{\dots}\oplus\mathbf{r_{i-1}}$. The policy model then decomposes the problem into a subproblem $\mathbf{d_{i}}$, following the policy $\pi(\mathbf{d_{i}}\mid\mathbf{q},\mathbf{s_{i-1}})$. Based on the subproblem $\mathbf{d_{i}}$ and the retrieved document, the intermediate result $\mathbf{r_{i}}$ is then generated and appended to the reasoning state repeatedly. Each step encodes an $(\mathbf{s_{i-1}},\mathbf{a_{i}})$ pair, where $\mathbf{s_{i-1}}$ represents the state, and $\mathbf{a_{i}}$ is a set $\{\mathbf{d_{i}},\mathbf{r_{i}}\}$ that implicitly reflects the step’s two actions, with $\mathbf{d_{i}}$ being the outcome corresponding to the decomposition and $\mathbf{r_{i}}$ to the retrieval-then-reasoning. A sequence of coherent steps, extending until the endpoint, collectively forms a complete trajectory.
@@ 2.2 Monte Carlo Tree Search
The MCTS algorithm expands a single trajectory into a search tree structure. The whole process begins with the original problem as the root node, followed by iterative searches consisting of selection, expansion, simulation, and back-propagation. The four phases are detailed as follows:
Selection.  Starting from the root node and traversing the existing tree structure, the algorithm selects the optimal child node in preparation for the next expansion phase. To balance exploration and exploitation, the well-known Upper Confidence Bounds (UCT) (Kocsis and Szepesvári, 2006) is employed in the selection process, formulated as:
@@
where $N(\mathbf{s},\mathbf{a})$ and $N(Pa(\mathbf{s}))$ represent the visit counts of the current node and its parent node in previous searches, respectively. The initial value of $Q(\mathbf{s},\mathbf{a})$ is calculated by the Risk-Value function (detailed in § 2.3) and is subsequently updated during the back-propagation phase.
@@
Back-propagation.  The back-propagation phase updates the values of all nodes along the selected reasoning branch. This process follows a bottom-up manner, where the parent node’s value is determined by the values and visit counts of its child nodes. The mathematical formulation is as follows:
@@ 2.3 Risk Assessment
In this section, we delve into the Risk-Value function, which risk-assesses the reasoning states to guide the tree search process. To begin with, for a composite problem $\mathbf{q}$, we treat its decomposition and retrieval-then-reasoning as a statistical decision of a probabilistic process (Zhai and Lafferty, 2006; Lafferty and Zhai, 2001). Specifically, given a set of decomposed subproblems $D=\{\mathbf{d^{1}},\mathbf{d^{2}},\mathbf{\dots},\mathbf{d^{k}}\}$ and the corresponding set of intermediate results $R=\{\mathbf{r^{1}},\mathbf{r^{2}},\mathbf{\dots},\mathbf{r^{k}}\}$ [^3] In this notation, the subscript of the symbol denotes the sequence number of the reasoning step, while the superscript indicates the identifier for different reasoning perspectives., the quality of a node state can be evaluated using a relevance score $p(\mathbf{r}\mid\mathbf{q}),\mathbf{r}\in R$ (Sachan et al., 2022). We substitute the “problem generation likelihood” (Zhai and Lafferty, 2001; Ponte and Croft, 1998) as an alternative to the relevance score after applying the Bayes’ rule:
$\log(\mathbf{r}\mid\mathbf{q})=\log p(\mathbf{q}\mid\mathbf{r})+\log p(\mathbf{r})+c,$
where $p(\mathbf{r})$ is the prior belief that $\mathbf{r}$ is relevant to any problem and is assumed to be uniform in this case. We can also drop $c$ since it is the intermediate results-independent constant. The formula is then reduced to:
$\log(\mathbf{r}\mid\mathbf{q})\propto\log p(\mathbf{q}\mid\mathbf{r}),\forall\mathbf{r}\in R,$
where $p(\mathbf{q}\mid\mathbf{r})$ captures how well the intermediate results $\mathbf{r}$ fits the particular problem $\mathbf{q}$. We utilize the policy model to compute the average log-likelihood of generating the original problem tokens in order to estimate $\log p(\mathbf{q}\mid\mathbf{r})$ (Sachan et al., 2022; Yuan et al., 2024), and define the expected risk of a node $(\mathbf{s},\mathbf{a})$ pointing to $\mathbf{r}$ as follows:
@@
where $\alpha,\beta$ is the translation and scaling factors.
@@ 3.1 Setup
We use HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MusiQue (Trivedi et al., 2022) as the test set. These datasets span a wide range of topics, necessitating the retrieval and reasoning over multiple supporting documents. To balance computational efficiency and evaluation robustness, we conduct experiments on a subset of 200 randomly selected questions (Jiang et al., 2024; Feng et al., 2025). More details are available in § B.1.
@@
All baselines are incorporated with RAG. For prompt-based baselines, we compare ARise with Query2Doc (Wang et al., 2023a), Self-Ask (Press et al., 2023), Verify-and-Edit (Zhao et al., 2023a) and Auto-RAG (Yu et al., 2024b). For search-based baselines, we use Self-Consistency (SC) (Wang et al., 2023b) and Retrieval-augmented-thought-tree (RATT) (Zhang et al., 2024b). We choose EM accuracy and F1 score as the evaluation metrics. The prediction is correct if the ground truth answer is exactly contained (Jiang et al., 2024; Feng et al., 2025). More details are available in § B.2.
@@ 3.2 Main Results
Table 1 presents the comprehensive experimental results on ARise and diverse baselines. Specifically, on the Qwen2.5-14B-Instruct model, ARise outperforms across all benchmarks, achieving an absolute improvement of 19.83% in EM over the vanilla RAG method, 13.29% over prompt-based baselines, and 15.5% over search-based baselines. ARise maintains robust performance on the Qwen2.5-7B-Instruct model with an absolute improvement of 13.67% in EM over the vanilla RAG method and overall surpasses various baselines. We observed that ARise performs slightly worse on Llama models. To further analyze this, we identify another interesting phenomenon: Auto-RAG, which adopts the same interleaved decomposition and retrieval paradigm as ARise, also exhibits a decline on Llama. This phenomenon suggests that Llama model may not be well-suited for iterative problem decompositions. In contrast, continuous-step reasoning approaches such as CoT and tree-of-thoughts show better results. Nevertheless, ARise still maintains a notable F1 advantage on Llama, indicating its effectiveness in selecting more promising paths.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities and are receiving increasing attention to enhance their reasoning through scaling test-time compute. However, their application in open-ended, knowledge-intensive, complex reasoning scenarios is still limited. Reasoning-oriented methods struggle to generalize to open-ended scenarios due to implicit assumptions of complete world knowledge. Meanwhile, knowledge-augmented reasoning (KAR) methods fail to address two core challenges: 1) error propagation, where errors in early steps cascade through the chain, and 2) verification bottleneck, where the explore–exploit trade-off arises in multi-branch decision processes. To overcome these limitations, we introduce ARise, a novel framework that integrates risk assessment of intermediate reasoning states with dynamic retrieval-augmented generation (RAG) within a Monte Carlo tree search paradigm. This approach enables effective construction and optimization of reasoning plans across multiple maintained hypothesis branches. Experimental results show that ARise significantly outperforms the state-of-the-art KAR methods by up to 23.10%, and the latest RAG-equipped large reasoning models by up to 25.37%. Our project page is at https://opencausalab.github.io/ARise.
@@
Yize Zhang1,2,3thanks: Equal contribution. Tianshu Wang5,7,811footnotemark: 1thanks: Work done during an internship at SenseTime. Sirui Chen1,6 Kun Wang4 Xingyu Zeng4 Hongyu Lin5 Xianpei Han5 Le Sun5 Chaochao Lu1,2thanks: Corresponding author. 1Shanghai AI Laboratory  2Shanghai Innovation Institute  3Shanghai Jiao Tong University 4SenseTime  5Institute of Software, Chinese Academy of Sciences  6Tongji University 7Hangzhou Institute for Advanced Study, 8University of Chinese Academy of Sciences ez220523@sjtu.edu.cn, tianshu2020@iscas.ac.cn, luchaochao@pjlab.org.cn
@@ 1 Introduction
Figure 1: Error propagation and verification bottleneck. Prior knowledge-augmented reasoning works fail to address two core challenges: 1) error propagation, where errors in early steps cascade through the chain, and 2) verification bottleneck, where the explore–exploit trade-off arises in multi-branch decision processes.
@@
Meanwhile, current knowledge-augmented reasoning (KAR) methods, as illustrated in Figure 1, are hindered by error propagation and verification bottleneck, which undermine reasoning reliability. To acquire knowledge for reasoning, retrieval-augmented generation (RAG) has been shown to be an effective way of dynamically retrieving documents as intermediate results Lewis et al. (2020); Liu et al. (2024). Prompt-based methods further extend KAR through chain-of-thought (CoT) prompting, which decomposes complex reasoning into sub-steps and iteratively retrieves relevant knowledge as reasoning proceeds Zhao et al. (2023a); Yu et al. (2024b); Li et al. (2024). However, this approach is plagued by error propagation, where errors in early steps can cascade through the chain. While search-based methods can mitigate error propagation by maintaining multiple hypothesis branches, verification bottleneck limits the effective explore–exploit trade-off in KAR. Existing verification solutions remain unsatisfactory as they rely on error-prone self-verification Stechly et al. (2024); Wang et al. (2023b); Zhang et al. (2024b), or on specific verifier training Setlur et al. (2024); Zhang et al. (2024a).
To overcome these limitations, we present a novel framework, ARise, towards knowledge-Augmented Reasoning via risk-adaptive search. As shown in Figure 2, ARise consists of three components: reasoning state generation, Monte Carlo tree search (MCTS), and risk assessment. Specifically, ARise iteratively refines reasoning steps through decomposition, retrieval-then-reasoning to provide fine-grained knowledge for LLMs (§ 2.1). MCTS treats each step as a node in the search tree, expanding linear reasoning to mitigate error propagation by enabling focused exploration of promising reasoning states and allowing backtracking when necessary (§ 2.2). Risk assessment leverages Bayesian risk minimization to evaluate the uncertainty of each state, dynamically balancing explore–exploit trade-off to guide the search towards both reliable and novel reasoning directions (§ 2.3). In this way, ARise enables robust and efficient complex reasoning by combining structured decomposition, knowledge retrieval, and risk-adaptive exploration in a unified framework.
We conducted comprehensive experiments with multiple LLMs on three challenging multi-hop question answering (QA) benchmarks that require complex reasoning and knowledge integration. Experimental results demonstrate that ARise significantly outperforms the state-of-the-art (SOTA) KAR methods, with an average of 23.10% and 15.52% improvement in accuracy and F1. In addition, when compared to the latest LRMs DeepSeek-AI et al. (2025) equipped with RAG, ARise also improve the average accuracy and F1 of 4.04% and 25.37%. These results verify the effectiveness of ARise for open-ended, knowledge-intensive, complex reasoning tasks.
@@
  • We propose a knowledge-augmented framework for open-ended complex reasoning and design a risk-adaptive MCTS algorithm to balance explore-exploit trade-off for reasoning.
  • We conduct comprehensive experiments to verify the effectiveness of ARise and to demonstrate that it outperforms the SOTA KAR methods and the latest LRMs equipped with RAG.
  • We provide empirical insights that 1) search-based wide reasoning can explore more solutions than learning-based deep reasoning, and 2) ARise progressively approaches optimal performance through model size scaling.
@@ 2 The ARise Method
Figure 2: Pipeline of ARise. ARise iteratively refines reasoning steps through decomposition, retrieval-then-reasoning, providing fine-grained knowledge for LLMs (§ 2.1). MCTS treats each step as a node in the search tree, expanding linear reasoning to mitigate error propagation by enabling exploration of reasoning paths and allowing backtracking when necessary (§ 2.2). Risk assessment leverages Bayesian risk minimization to evaluate the quality of each state, dynamically optimizing action strategies to guide the search towards promising directions (§ 2.3).
@@
  • Monte Carlo Tree Search: MCTS transforms a sequence of interconnected nodes into a tree structure. Each node can undergo a simulated rollout under guidance. The local value incorporating future reward can be updated without incurring the cost of taking an actual forward step.
  • Risk Assessment: We designed the Risk-Value function to assess the risks of the intermediate reasoning states at each node. The policy model is capable of dynamically formulating and adjusting action strategies based on the actual risk associated with each branch.
@@ 2.1 Reasoning State Generation
To define steps more clearly and granularly, we prompt LLMs to perform problem decomposition and retrieval-then-reasoning in an interleaved manner. These two consecutive actions together constitute a single step. Intermediate results at each step are continuously appended to the entire reasoning state and serve as new inputs for subsequent steps, progressively approaching the final solution to complex tasks. This approach, where intermediate subtasks and their labels are concatenated to the original task’s input to form a new input sequence, is widely applied in compounded tasks (Wies et al., 2023; Rajani et al., 2019; Cobbe et al., 2021). Specifically, at the $i^{\text{th}}$ step [^2] We use the bold notation for vectors and non-bold notation for scalars., the input comprises the original problem $\mathbf{q}$ and the intermediate results ${\mathbf{r_{1}},\mathbf{r_{2}},\mathbf{\dots},\mathbf{r_{i-1}}}$ from previous steps, with the latter forming the reasoning state $\mathbf{s_{i-1}}=\mathbf{r_{1}}\oplus\mathbf{r_{2}}\oplus\mathbf{\dots}\oplus\mathbf{r_{i-1}}$. The policy model then decomposes the problem into a subproblem $\mathbf{d_{i}}$, following the policy $\pi(\mathbf{d_{i}}\mid\mathbf{q},\mathbf{s_{i-1}})$. Based on the subproblem $\mathbf{d_{i}}$ and the retrieved documents, the intermediate result $\mathbf{r_{i}}$ is then generated and appended to the reasoning state repeatedly. Each step encodes an $(\mathbf{s_{i-1}},\mathbf{a_{i}})$ pair, where $\mathbf{s_{i-1}}$ represents the state, and $\mathbf{a_{i}}$ is a set $\{\mathbf{d_{i}},\mathbf{r_{i}}\}$ that implicitly reflects the step’s two actions, with $\mathbf{d_{i}}$ being the outcome corresponding to the decomposition and $\mathbf{r_{i}}$ to the retrieval-then-reasoning. A sequence of coherent steps, extending until the endpoint, collectively forms a complete trajectory.
@@ 2.2 Monte Carlo Tree Search
The MCTS algorithm expands a single trajectory into a search tree structure. The whole process begins with the original problem as the root node, followed by iterative searches consisting of selection, expansion, simulation, and backpropagation. The four phases are detailed as follows:
Selection.  Starting from the root node and traversing the existing tree structure, the algorithm selects the optimal child node in preparation for the next expansion phase. To balance exploration and exploitation, the well-known Upper Confidence Bounds (UCT) (Kocsis and Szepesvári, 2006) is used in the selection process, formulated as:
@@
where $N(\mathbf{s},\mathbf{a})$ and $N(Pa(\mathbf{s}))$ represent the visit counts of the current node and its parent node in previous searches, respectively. The initial value of $Q(\mathbf{s},\mathbf{a})$ is calculated by the Risk-Value function (detailed in § 2.3) and is subsequently updated during the backpropagation phase.
@@
Backpropagation.  The backpropagation phase updates the values of all nodes along the selected reasoning branch. This process follows a bottom-up manner, where the parent node’s value is determined by the values and visit counts of its child nodes. The mathematical formulation is as follows:
@@ 2.3 Risk Assessment
In this section, we delve into the Risk-Value function, which assesses the risks of reasoning states to guide the tree-search process. To begin with, for a composite problem $\mathbf{q}$, we treat its decomposition and retrieval-then-reasoning as a statistical decision of a probabilistic process (Zhai and Lafferty, 2006; Lafferty and Zhai, 2001). Specifically, given a set of decomposed subproblems $D=\{\mathbf{d^{1}},\mathbf{d^{2}},\mathbf{\dots},\mathbf{d^{k}}\}$ and the corresponding set of intermediate results $R=\{\mathbf{r^{1}},\mathbf{r^{2}},\mathbf{\dots},\mathbf{r^{k}}\}$ [^3] In this notation, the subscript of the symbol denotes the sequence number of the reasoning step, while the superscript indicates the identifier for different reasoning perspectives., the quality of a node state can be evaluated using a relevance score $p(\mathbf{r}\mid\mathbf{q}),\mathbf{r}\in R$ (Sachan et al., 2022). We substitute the “problem generation likelihood” (Zhai and Lafferty, 2001; Ponte and Croft, 1998) as an alternative to the relevance score after applying the Bayes’ rule:
$\log p(\mathbf{r}\mid\mathbf{q})=\log p(\mathbf{q}\mid\mathbf{r})+\log p(\mathbf{r})+c,$
where $p(\mathbf{r})$ is the prior belief that $\mathbf{r}$ is relevant to any problem and is assumed to be uniform in this case. We can also drop $c$ since it is the intermediate result-independent constant. The formula is then reduced to:
$\log p(\mathbf{r}\mid\mathbf{q})\propto\log p(\mathbf{q}\mid\mathbf{r}),\forall\mathbf{r}\in R,$
where $p(\mathbf{q}\mid\mathbf{r})$ captures how well the intermediate results $\mathbf{r}$ fit the particular problem $\mathbf{q}$. We utilize the policy model to compute the average log-likelihood of generating the original problem tokens in order to estimate $\log p(\mathbf{q}\mid\mathbf{r})$ (Sachan et al., 2022; Yuan et al., 2024), and define the expected risk of a node $(\mathbf{s},\mathbf{a})$ pointing to $\mathbf{r}$ as follows:
@@
where $\alpha,\beta$ are the translation and scaling factors.
@@ 3.1 Setup
We use HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MusiQue (Trivedi et al., 2022) as test sets. These datasets span a wide range of topics, necessitating the retrieval and reasoning over multiple supporting documents. To balance computational efficiency and evaluation robustness, we conduct experiments on a subset of 200 randomly selected questions (Jiang et al., 2024; Feng et al., 2025). More details are available in § B.1.
@@
All baselines are incorporated with RAG. For prompt-based baselines, we compare ARise with Query2Doc (Wang et al., 2023a), Self-Ask (Press et al., 2023), Verify-and-Edit (Zhao et al., 2023a) and Auto-RAG (Yu et al., 2024b). For search-based baselines, we use Self-Consistency (SC) (Wang et al., 2023b) and Retrieval-augmented-thought-tree (RATT) (Zhang et al., 2024b). We choose EM accuracy and F1 score as our evaluation metrics. The prediction is correct if the ground truth answer is exactly contained (Jiang et al., 2024; Feng et al., 2025). More details are available in § B.2.
@@ 3.2 Main Results
Table 1 presents the comprehensive experimental results on ARise and various baselines. Specifically, on the Qwen2.5-14B-Instruct model, ARise outperforms across all benchmarks, achieving an absolute improvement of 19.83% in EM over the vanilla RAG method, 13.29% over prompt-based baselines, and 15.5% over search-based baselines. ARise maintains robust performance on the Qwen2.5-7B-Instruct model with an absolute improvement of 13.67% in EM over the vanilla RAG method and overall surpasses various baselines. We observed that ARise performs slightly worse on Llama models. To further analyze this, we identify another interesting phenomenon: Auto-RAG, which adopts the same interleaved decomposition and retrieval paradigm as ARise, also exhibits a decline on Llama. This phenomenon suggests that the Llama model may not be well-suited for iterative problem decompositions. In contrast, continuous-step reasoning methods such as CoT and tree-of-thoughts show better results. Nevertheless, ARise still maintains a notable F1 advantage on Llama, indicating its effectiveness in selecting more promising paths.